OpenStudy (anonymous):

Find all the real and imaginary solution to the equation 6x^4-5x^3+7x^2-5x+1=0

OpenStudy (anonymous):

Help pls

OpenStudy (anonymous):

man they did not make this easy did they?

OpenStudy (anonymous):

possible rational roots are \[\pm1,\pm\frac{1}{2},\pm\frac{1}{3},\pm\frac{1}{6}\]

OpenStudy (anonymous):

turn out the ones that work are \(\frac{1}{2}\) and \(\frac{1}{3}\)

OpenStudy (anonymous):

Then after wat do I do to find the solution